# BatchLLM: Optimizing Large Batched LLM Inference with Global Prefix Sharing and Throughput-oriented Token Batching

### v2

@@ 7. Related Work
Prefix sharing systems. The vLLM (Kwon et al., 2023) proposes the PagedAttention to manage the KV memory in blocks, which enables to reuse the KV context both intra and inter requests in memory block level. Prompt Cache (Gim et al., 2024) and SGLang (Zheng et al., 2023) propose the domain specific language (DSL) to define the shared prefix of prompt. Besides, SGLang proposes the radix tree based KV cache management with LRU policy for KV memory eviction. RAGCache (Jin et al., 2024b) studies the common KV caching for RAG optimization. Some recent works (Gao et al., 2024; Jin et al., 2024a; Qin et al., 2024; Hu et al., 2024; OpenAI, 2024b) study the distributed request scheduling with the consideration of prompt prefix reuse. None of these works identify the common prefix ahead-of-time for the large batch and enable the explicit reusing. Instead, BatchLLM manages the prefix reusing explicitly and can reuse the KV context the best for the large batched scenario.
@@
Token batching and serving optimization. Orca (Yu et al., 2022) proposes the continuous batching method for transformer models with the per-iteration token batching of auto-regressive models (Gao et al., 2018). FastGen (Holmes et al., 2024) and SARATHI (Agrawal et al., 2023) study the mix of of prefill chunks and decoding tokens into the same token-batch to increase the compute intensity of the batch. FlexGen (Sheng et al., 2023) proposes the swizzled token computation and offloading to support the LLM inference with limited GPU memory. Some works (Fu et al., 2024; Sheng et al., 2024) study the request scheduling to achieve better SLO for online LLM serving. BatchLLM differs from these works as it targets the large batched inference, leveraging the static information to reorder the requests and using the memory-aware scheduling to enlarge the size of token-batch. The vLLM (Kwon et al., 2023) uses the multi-step scheduling method to reduce the token-batch scheduling overhead, which is orthogonal to BatchLLM. Another orthogonal dimension to improve the token-batch performance is the pruning and quantization (Xia et al., 2023; Frantar et al., 2022; Lin et al., 2024), which can be applied concurrently with the optimizations in this paper.
@@
Attention kernel optimization. Memory-efficient Attention (Rabe and Staats, 2021) and FlashAttention (Dao et al., 2022) fuse the Attention operators into a single kernel to boost the performance, with Online Softmax (Milakov and Gimelshein, 2018) to tile the Attention computation. The recent prefix-shared Attention optimization (ChunkAttention (Ye et al., 2024b), RelayAttention (Zhu et al., 2024), Hydragen (Juravsky et al., 2024), and Cascade Inference (Ye et al., 2024a)) compute the attention on the prefix and other part separately in different kernels, for which the former is transformed from matrix-vector multiplication between Q and K/V into MatMul, and reduce the results of different parts according to the Online Softmax algorithm. Different from these works, BatchLLM fuses the Attention computation of different parts into the same kernel to reduce tail latency and kernel launch overhead. The horizontal fusion has been proposed and used in the existing machine learning optimizers (Pan et al., 2023, 2024; Li et al., 2022). BatchLLM borrows this idea and applies it in the prefix-shared Attention.
8. Conclusion
@@
This paper presents BatchLLM, the holistic optimization techniques for large batched LLM-based information processing and management tasks. It identifies the limitations of existing methods, and optimizes the tasks with the global information of the large batch. It explicitly extracts the common prefix globally to avoid prefix’s early eviction problem, and simplifies the prefix pattern by enlarging the first-level prefix with a DP algorithm on the tree to reduce the overhead of scheduling and Attention computation. It schedules the requests at the granularity of prefix-sharing groups, which enables the global prefix sharing the best and shrinks the lifetime of prefix’s KV memory. It proposes the request reordering and memory-centric token batching method to better mix the prefill chunks into the decoding token-batches and thus better saturates the GPU. Finally, it presents the horizontal fused prefix-shared Attention kernel to reduce the tail effect and kernel launch overhead. Extensive evaluation shows that BatchLLM outperforms vLLM and SGLang by 1.3×1.3$\times$1.3 × to 10.8×10.8$\times$10.8 × on a set of microbenchmarks and a typical industry workload under different hardware environments.

### v3

@@ 5 Related Work
Prefix sharing systems. The vLLM Kwon et al. (2023) proposes the PagedAttention to manage the KV memory in blocks, which enables to reuse the KV context both intra and inter requests in memory block level. Prompt Cache Gim et al. (2024) and SGLang Zheng et al. (2023) propose the domain specific language (DSL) to define the shared prefix of prompt. Besides, SGLang proposes the radix tree based KV cache management with LRU policy for KV memory eviction. RAGCache Jin et al. (2024a) studies the common KV caching for RAG optimization. Some recent works Gao et al. (2024); Jin et al. (2024b); Qin et al. (2024); Hu et al. (2024); OpenAI (Cited 2024b) study the distributed request scheduling with the consideration of prompt prefix reuse. None of these works identify the common prefix ahead-of-time and enable the explicit reusing. Instead, BatchLLM manages the prefix reusing explicitly and can reuse the KV context the best for the large batched scenario.
@@
Token batching and serving optimization. Orca Yu et al. (2022) proposes the continuous batching method for transformer models with the per-iteration token batching of auto-regressive models Gao et al. (2018). FastGen Holmes et al. (2024) and SARATHI Agrawal et al. (2023) study the mix of prefill chunks and decoding tokens into the same token-batch to increase the compute intensity of the batch. FlexGen Sheng et al. (2023) proposes the swizzled token computation and offloading to support the LLM inference with limited GPU memory. Some works Fu et al. (2024); Sheng et al. (2024) study the request scheduling to achieve better SLO for online LLM serving. BatchLLM differs from these works as it targets the large batched inference, leveraging the static information to reorder the requests and using the memory-aware scheduling to enlarge the size of token-batch. The vLLM Kwon et al. (2023) uses the multi-step scheduling method to reduce the token-batch scheduling overhead, which is orthogonal to BatchLLM. Another orthogonal dimension to improve the token-batch performance is the pruning and quantization Xia et al. (2023); Frantar et al. (2022); Lin et al. (2024), which can be applied concurrently with the optimizations in this paper.
@@
Attention kernel optimization. Memory-efficient Attention Rabe and Staats (2021) and FlashAttention Dao et al. (2022) fuse the Attention operators into a single kernel to boost the performance, with Online Softmax Milakov and Gimelshein (2018) to tile the Attention computation. The recent prefix-shared Attention optimization (ChunkAttention Ye et al. (2024a), RelayAttention Zhu et al. (2024), Hydragen Juravsky et al. (2024), and Cascade Inference Ye et al. (2024b)) compute the attention on the prefix and other part separately in different kernels, for which the former is transformed from matrix-vector multiplication between Q and K/V into MatMul, and reduce the results of different parts through Online Softmax. Different from these works, BatchLLM fuses the Attention computation of different parts into the same kernel to reduce tail latency and kernel launch overhead. The horizontal fusion has been proposed and used in the existing machine learning optimizers Pan et al. (2023; 2024); Li et al. (2022). BatchLLM borrows this idea and applies it in the prefix-shared Attention.
6 Conclusion
@@
This paper presents BatchLLM, the holistic optimization techniques for large batched LLM-based information processing and management tasks. It identifies the limitations of existing methods, and optimizes the tasks with the global information of the large batch. It explicitly extracts the common prefix globally to avoid prefix’s early eviction problem, and simplifies the prefix pattern by enlarging the first-level prefix with a DP algorithm on the tree to reduce the overhead of scheduling and Attention computation. It schedules the requests at the granularity of prefix-sharing groups, which enables the global prefix sharing the best and shrinks the lifetime of prefix’s KV memory. It proposes the request reordering and memory-centric token batching method to better mix the prefill chunks into the decoding token-batches and thus better saturates the GPU. Finally, it presents the horizontal fused prefix-shared Attention kernel to reduce the tail effect and kernel launch overhead. Extensive evaluation shows that BatchLLM outperforms vLLM and SGLang by $1.3\times$ to $10.8\times$ on a set of microbenchmarks and a typical industry workload under different hardware environments.